Circular No. 6980

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998de IN NGC 252 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum of SN 1998de (cf. IAUC 6977) was obtained with the Tillinghast 1.5-m telescope on July 25.4 UT by M. Calkins. A broad absorption feature of Si II is seen at 618.5 nm, implying that this is a type-Ia supernova. Correcting for the host-galaxy redshift of 5023 km/s (from the NASA Extragalactic Database), the photospheric expansion velocity is 13 300 km/s. The blue end of the spectrum shows Ti II absorption, characteristic of the sub-luminous SNe 1991bg and 1986G, and the very strong Si II at 580 nm also suggests that this will be a fastdeclining type-Ia event.

SUPERNOVA 1998dh IN NGC 7541 Garnavich et al. also write: “A spectrum of SN 1998dh (cf. IAUC 6978)

was obtained with the Tillinghast 1.5-m telescope on July 26.4 UT by P. Berlind. The spectrum is clearly that of a normal type-Ia supernova before maximum. Broad Si II is observed at 613.21 nm, implying a photospheric velocity of 13 400 km/s (assuming a redshift for the host galaxy of 2678 km/s from the NASA Extragalactic Database). Interstellar absorption lines of Na I are seen from our galaxy with an equivalent width (EW) of 0.1 nm and from NGC 7541 with an EW of 0.04 nm. SN 1998dh should reach a visual magnitude of ∼ 13.8.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: June 20.37 UT, 11.7 (R. Keen, Mt.

Thorodin, CO, 0.32-m reflector); 28.44, 11.4 (N. Biver, Oahu, HI, 0.26-m reflector); July 2.90, 10.8 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 16.95, 11.6 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 19.16, 10.9 (J. Bortle, Stormville, NY, 0.41-m reflector); 22.98, 11.3 (A. Baransky, Kiev, Ukraine, 0.20-m reflector); 26.90, 10.9 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector).

SGR 1627–41

Corrigendum. IAUC 6962 titled this object incorrectly.

1998 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green